OpenStudy (anonymous):

LOG(10-4X)=LOG(10-3X)

OpenStudy (anonymous):

\[\log_{10} (10-4x) = \log_{10}(10-3x)\] I think the only solution to this problem is when x = 0... :/

OpenStudy (anonymous):

LOG(10-4X)=LOG(10-3X) cancel out the LOG from both sides we have; (10-4X)=(10-3X) 10 -4X = 10-3X 10-10 = 4X-3X 0= X OR X=0 IS THE SOLUTION OF THE GIVEN PROBLEM. @MATH5